# VIVAT: Virtuous Improving VAE Training through Artifact Mitigation

### v1

@@ 6.3 Text-to-image generation
To assess the integration of our VAE with latent diffusion pipeline, we train a Diffusion Transformer (DiT), which is similar to LI-DiT model [29]. The architecture consists of two principal components:
  • Token Refiner that processes textual embeddings through self-attention and feed-forward layers with Adaptive Layer Normalization (AdaNorm) [52] for temporal conditioning;
@@
For comparative evaluation between VIVAT and Flux VAE [26], we trained a 2 billion parameter DiT model for 130 thousand steps on $256\times 256$ resolution images with a batch size of 1024. Generation quality was quantified using the CLIP-score metric [17] to measure text-image alignment. Figure 6 shows that our method allows for a stable improvement in quality during the training process and surpasses the pipeline with the same diffusion transformer, but using Flux VAE. Details of the architecture and training of the diffusion transformer can be found in Appendix A.3.
@@ A.1 VAE Blocks
Figure 7: The ResNet Block for the VAE Encoder or Decoder. We use the Group Normalization [51] for the Encoder and Spatially Conditional Normalization [58] for the Decoder.
Figure 8: The Self-Attention Block for the VAE Encoder or Decoder. We use the Group Normalization [51] for the Encoder and Spatially Conditional Normalization [58] for the Decoder.

### v2

@@ 6.3 Text-to-image generation
To assess the integration of our VAE with latent diffusion pipeline, we train a Diffusion Transformer (DiT), which is similar to LI-DiT model [ma2024exploring]. The architecture consists of two principal components:
  • Token Refiner that processes textual embeddings through self-attention and feed-forward layers with Adaptive Layer Normalization (AdaNorm) [xu2019understandingimprovinglayernormalization] for temporal conditioning;
@@
For comparative evaluation between VIVAT and Flux VAE [flux2024], we trained a 2 billion parameter DiT model for 130 thousand steps on $256\times 256$ resolution images with a batch size of 1024. Generation quality was quantified using the CLIP-score metric [Hessel2021CLIPScoreAR] to measure text-image alignment. Figure 6 shows that our method allows for a stable improvement in quality during the training process and surpasses the pipeline with the same diffusion transformer, but using Flux VAE. Details of the architecture and training of the diffusion transformer can be found in Appendix A.3.
@@ A.1 VAE Blocks
Figure 7: The ResNet Block for the VAE Encoder or Decoder. We use the Group Normalization [wu2018groupnormalization] for the Encoder and Spatially Conditional Normalization [zheng2022movq] for the Decoder.
Figure 8: The Self-Attention Block for the VAE Encoder or Decoder. We use the Group Normalization [wu2018groupnormalization] for the Encoder and Spatially Conditional Normalization [zheng2022movq] for the Decoder.